OpenStudy (anonymous):

Factor 16x^2 + 24x + 9. (16x + 3)(x + 3) (4x + 3)(4x - 3) (4x + 3)^2 i got b is that correct

OpenStudy (a1234):

(4x + 3)(4x - 3) 16x^2 - 12x + 12x - 9 16x^2 - 9 No, b is not it.

OpenStudy (anonymous):

k ill try it again

OpenStudy (anonymous):

does factor mean simplify ?

OpenStudy (anonymous):

(4x)^2+2*4x*3+3^2=(4x+3)^2

OpenStudy (a1234):

To factor the above problem, I think you need to find two numbers that have a sum of 24 and a product of 144.

OpenStudy (anonymous):

c is correct

OpenStudy (anonymous):

kk its 12

OpenStudy (anonymous):

it is a perfect square

OpenStudy (anonymous):

dont give the awnser

OpenStudy (a1234):

Yup, 12.

OpenStudy (anonymous):

it is a perfect square man open your eyes @a1234

OpenStudy (anonymous):

why to go behind factor

OpenStudy (anonymous):

16*9=144 12+12=24 12*12=144 write 24x=12x+12x and make factors.

OpenStudy (anonymous):

and make it long and calculative

OpenStudy (a1234):

What do you mean? He's asking how to factor the problem. Suppose you don't have the answer choices in front of you, then how do you do it? @g17

OpenStudy (a1234):

It is not always a perfect square.

OpenStudy (anonymous):

(a+b)^2 hope u remember it

OpenStudy (a1234):

Okay...suppose you have the following problem. (13x^2 + 17x + 4) Then how do you do it? It is not a perfect square. I was just telling dirtydan667 what steps to use.

OpenStudy (gorv):

then u should have taken a common approch

OpenStudy (gorv):

it was ur mistake that u took a perfect square as example

OpenStudy (a1234):

It was an example...it can be used to any problem. Even behind a perfect square, there are details you have to know. If you did a perfect square without knowing why you perform each step, it is senseless.